OpenStudy (anonymous):

use the intercepts to graph the equation 2x+5y=10

hero (hero):

2(0) + 5y = 10 5y = 10 y =2 2x + 5(0) = 10 2x = 10 x = 5 Graph the points (0,2) and (5,0) on graph paper Find a straight edge and draw a line through the points Write 2x + 5y = 10 above it